OpenStudy (anonymous):

sin^4x + cosx^4x = 1 ~.~ someone help please =^= ??

OpenStudy (debbieg):

\[\sin^4x + cosx^4x = \sin^4x + (1-\sin^2x)^2\] expand the RHS and then let\[u=\sin^2x\] And see if you can't get a nice easy little quadratic to solve. :)

OpenStudy (anonymous):

uhm.... can't you explain? :< =.=

OpenStudy (debbieg):

You are supposed to solve the equation, right?

OpenStudy (anonymous):

right ...

OpenStudy (debbieg):

Wait - is that a typo? I just realized, you have (and I copied) \[\sin^4x + cosx^4x = 1\] I'm assuming it should be\[\sin^4x + \cos^4x = 1\]

OpenStudy (debbieg):

so \[ \cos^4x = (\cos^2x)^2\] and\[ (\cos^2x)^2=(1-\sin^2x)^2\] right?

OpenStudy (anonymous):

yup ! and ... :<

OpenStudy (anonymous):

kimmese did u understand the basic concept ?

OpenStudy (debbieg):

So we can re-write the equation:\[\sin^4x + (1-\sin^2x)^2 = ?\] Expand that, and you will have an equation all in sinx

OpenStudy (debbieg):

I should say, "we can rewrite the LHS of the equation...."

OpenStudy (anonymous):

uhm cos^2x + sin^2 x =1 right ? so sin^2x = 1 - cos^2x right ... :<

OpenStudy (debbieg):

Yes, sorry. That's the fundamental ID, I was assuming that didn't need to be stated. :)

OpenStudy (anonymous):

:> =.= now i got it =.= ! thanks =.= !!

OpenStudy (debbieg):

you're welcome :)

OpenStudy (tkhunny):

I'm kind of fond of a slightly different method. Take \(\sin^{2}(x) + \cos^{2}(x) = 1\) Everything is positive, so there is no harm in squaring. \(\sin^{4}(x) + 2\sin^{2}(x)\cos^{2}(x) + \cos^{4}(x) = 1\) From the original equation, this gives: \(2\sin^{2}(x)\cos^{2}(x) = 0\)

OpenStudy (anonymous):

thank you =)